# The Role of Model Confidence on Bias Effects in Measured Uncertainties

### v1

@@ 6 Conclusion
While entropy decomposes into epistemic and aleatoric components, our findings show that lower model confidence amplifies bias effects on measured uncertainties, with a greater amplification observed on epistemic than on aleatoric uncertainty. Moreover, while model confidence does not significantly affect the direction of measured aleatoric changes under bias, lower model confidence is associated with greater underestimation of epistemic uncertainty (i.e. overconfidence) under bias.
Future work may leverage the distinct effects of bias on these two types of uncertainty across varying confidence levels to develop more advanced techniques for disentangling them.
@@ A.5 More Empirical Results
More specifically, the effects of bias, measured by changes in measured uncertainties, are more pronounced when model confidence is lower; in other words, when debiased epistemic entropy is higher. This is evidenced by consistently positive and statistically significant coefficients for changes in measured epistemic uncertainty due to biases in GPT-4o. Qwen2-VL follows the same pattern, with exceptions for Label bias. For aleatoric uncertainty, GPT-4o also shows predominantly positive coefficients, whereas Qwen2-VL exhibits inconsistent coefficient directions with much smaller values, as indicated by Epi./Ale. ratios greater than one-except for the same two exceptions, and nonsignificant p-values. These results are detailed in Table 8.
Lower model confidence is more strongly associated with greater underestimation of measured epistemic uncertainty, whereas it has no significant effect on the direction of changes in measured aleatoric uncertainty. This is supported by the consistently positive and largely significant coefficients for the decrease in measured epistemic uncertainty, while the coefficients for the decrease in measured aleatoric uncertainty are predominantly insignificant except the same two Qwen2-VL cases.

### v2

@@ 6 Conclusion
While uncertainty decomposes into epistemic and aleatoric components, our findings show that lower model confidence amplifies bias effects on measured uncertainties, with a greater amplification observed on epistemic than on aleatoric uncertainty. Moreover, while model confidence does not significantly affect the direction of bias-induced changes in measured aleatoric uncertainty, lower model confidence is associated with greater bias-induced underestimation of epistemic uncertainty (i.e. overconfidence).
Future work may leverage the distinct bias effects on these two types of uncertainty across varying confidence levels to develop more advanced techniques for disentangling them. In addition, our analysis could be extended to pure-text datasets once more challenging or less exposed multi-label benchmarks in LLMs pretraining become available.
@@ A.5 More Empirical Results
Bias-free Epistemic Entropy as Model Confidence
@@
More specifically, the effects of bias, measured by changes in measured uncertainties, are more pronounced when model confidence is lower; in other words, when debiased epistemic entropy is higher. This is evidenced by consistently positive and statistically significant coefficients for bias-induced changes in measured epistemic uncertainty on GPT-4o. Qwen2-VL follows the same pattern, with exceptions for Label bias. For aleatoric uncertainty, GPT-4o also shows predominantly positive coefficients, whereas Qwen2-VL exhibits inconsistent coefficient directions with much smaller values, as indicated by Epi./Ale. ratios greater than one with the same two exceptions, and nonsignificant p-values. These results are detailed in Table 9.
Lower model confidence is more strongly associated with greater underestimation of measured epistemic uncertainty, whereas it has no significant effect on the direction of bias-induced changes in measured aleatoric uncertainty. This is supported by the consistently positive and largely significant coefficients for the decrease in measured epistemic uncertainty, while the coefficients for the decrease in measured aleatoric uncertainty are predominantly insignificant except for the same two Qwen2-VL cases.
Extension to Claude-3.7-Sonnet.
Due to budget constraints, we randomly selected two biases for extended experiments on Claude-3.7-Sonnet with CREPE. Table 8 presents the results, which align with the findings observed on other models. More specifically, stronger bias effects on epistemic uncertainty estimation are associated with lower bias-free model confidence, as shown by negative and statistically significant coefficients. In contrast, the relationship for aleatoric uncertainty is weaker, as indicated by non-significant p-values and coefficients with mixed signs.
Bias Metrics Absolute Change Underestimation
Phrasing
Coefficients
-0.1579 0.0111 14.23 -0.1238 - -0.0184 6.73
P-value
ns
ns
Label
Coefficients
-0.5420 0.2012 2.69 -0.5681 0.2094 2.71
P-value
Table 8: This table reports the impact of phrasing and label bias on CREPE with Claude-3.7-Sonnet. Stronger bias effects on epistemic uncertainty estimation correspond to lower bias-free model confidence, as indicated by the negative coefficients with significant p-values. For aleatoric uncertainty, this relationship is weaker, reflected in the non-significant p-values and mixed coefficient signs.